OpenStudy (anonymous):

FIND THE VOLUME:The base is a right triangle with a leg of 12 in. and hypotenuse of 15 in.The height of the prism is 10 in.

OpenStudy (anonymous):

Since the hypotenuse is 15, by Pythagorean theorem (or triples), we get 9 as the other base. 9*12*1/2 gives us 54 for the triangle base. Since the prism is 10 in tall, all we do is multiply 54 by 10 to gives us 540 in^3. Hope this helps!

OpenStudy (anonymous):



OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

First find the 3rd side of the triangle. It is SQRT of 15^2 - 12^2 = SQRT(81) = 9. Area of that right triangle then will be: 1/2 * base * height = 1/2 * 12 * 9 = 6 * 9 = 54 square inches. Volume of that prism: Height times area = 54 * 10 = 540 cubic inches.